OpenStudy (anonymous):

2. Point slope form for m = 2, and point = (4,2) *

OpenStudy (anonymous):

y-2= 2(x-4)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you welcome

OpenStudy (anonymous):

Rewrite to have only integers. y = 1/3 x + 2